# CHAI: Command Hijacking against embodied AI

### v1

@@ VII Joint Semantic and Perceptual Optimization
The cross-entropy method defines a probability distribution over the search space $\Pi$, assigning higher probability mass to regions that yield better objective values. Let $\Omega$ denote the (unknown) optimal distribution. Our goal is to approximate $\Omega$ by a parametric distribution $p_{\alpha}$, with parameters $\alpha\in P$, that minimizes the divergence $\mathcal{KL}(\Omega\,,\,p_{\alpha})$. In the ideal case, $\mathcal{KL}(\Omega,p_{\alpha})=0$.
@@
By using the first-order optimality conditions, we can get the solution to Equation (5) in a closed form depending on the form of $p_{\alpha}$ and $\Omega$. In particular, we consider a piecewise uniform distribution and divide the space $\Pi$ into $m$ disjoint subsets $C_{1},...,C_{m}\subset\Pi$. Then, the jthj-th element of $\alpha$, denoted as $\alpha_{h,j}$, represents the probability associated with the maximum being in $C_{j}$. We then update $\alpha$ as,
[…]
with $\gamma_{i}=\frac{\Omega(\pi_{i})}{p_{\alpha_{i}}(\pi_{i})}$. For the complete derivation of this formula, refer to [33, 34].
@@ VIII-A Applications
We also implement a variation of DriveLM for experiments with real-world robotic vehicles where we deploy the attack on a printed surface.
@@
Tab. III presents the optimization variables for each application. For most applications, we optimize the sign letter background colors in RGB space and the prompt.
[…]
TABLE III: Optimization variables
@@ VIII-B Evaluation metrics
[…]
For this evaluation, we follow a similar approach to previous works [9], where we also evaluate the ASR in the scenario without an attack. This accounts for the stochasticity inherent to LVLMs; a query with the same image and figure may create different outputs.
@@ VIII-C Datasets
For each application we constructed two datasets: Known Images, used during attack optimization, and Transferability Images, held out to evaluate generalization (i.e., images that we do not show our optimizer). Figs. 68 illustrate this setup: each figure pairs an example from the Known Images set (left) with one from the Transferability Images set (right), highlighting the diverse conditions used to test CHAI on images unseen during optimization.
Details of the dataset construction, their diversity, and the labeling (including ASR errors in the benign case) can be found in Appendix A
Figure 6: Landing: Known (left) and Transferability (right).
Figure 7: DriveLM: Known (left) and Transferability (right).
Figure 8: Cloudtrack: Known (left) and Transferability (right).
@@ VIII-D Implementation details
LVLM: We test our attack against GPT-4o [36], a proprietary model, and InternVL2.5 8B [37], an open-weights model.
@@
Closed-loop implementation: We use AirSim [39], which is an open-source high-fidelity simulator for drones, to create a closed-loop implemenation of the landing application.
@@ IX Results
In this section we implement CHAI attacks for the applications described in Section VIII. We begin by evaluating CHAI on the Known images and benchmarking against SceneTAP, then assess transferability on held-out Transferability images. We then print physical signs of the optimized prompts and perform physical-world tests on a real robotic platform.
GPT InternVL Application No Attack SceneTAP CHAI No Attack SceneTAP CHAI Landing $0.00$ p m 0.00 $6.25$ p m 3.19 68.10 7.15 $15.50$ p m 8.96 $26.5$ p m 12.99 66.43  9.89 CloudTrack $0.00$ p m 0.00 $77.00$ p m 8.01 95.50 6.05 $14.50$ p m 8.87 $82.50$ p m 9.67 92.50  9.10 DriveLM $6.22$ p m 2.93 $55.67$ p m 6.22 81.78 5.00 $29.41$ p m 7.09 $47.67$ p m 12.85 54.74 13.07
TABLE IV: Attack Success Rate (ASR) in the Known Images Datasets.
@@ IX-A Attack success on Known Images
We implement CHAI and SceneTAP on Known Images. Tab. IV compares the ASR between both strategies. We draw the following conclusions.
@@ IX-B Attack Success on Transferability
We now study how our attack transfers to unknown scenarios. We apply the CHAI attacks that we obtain with the Known Images, apply them to the Transferability Images, and present the results in Tab. V. Notice that we cannot include SceneTAP in these results as SceneTAP is optimized only per image, and thus cannot generalize to previously unseen images.
@@
Application GPT InternVL No Attack CHAI No Attack CHAI Landing $0.00$ p m 0.00 71.38 6.34 $21.88$ p m 8.46 52.22 11.01 CloudTrack $0.00$ p m 0.00 95.50 7.59 $15.00$ p m 10.51 66.50  9.88 DriveLM $2.08$ p m 5.20 81.92 4.98 $33.08$ p m 15.90 51.25 15.12
TABLE V: ASR in the Transferability Images Datasets.
IX-C Real world deployment
Attack setup: We position a second robotic vehicle as an obstacle directly ahead of the victim vehicle so that, under benign conditions, the LVLM issues a stop command. The attacker places a printed adversarial prompt on or near the obstacle and aims to cause the LVLM to output a proceed command instead; an attack succeeds if this induced command causes the victim to move forward and collide with the obstacle.
We consider two attack scenarios, which we illustrate in Figs. 9(b)-9(c):
@@ IX-C Real world deployment
Tab. VI presents the ASR for both scenarios and LVLMs. When testing the case without an attack, we remove the objects where we place the CHAI signs. Consequently, the ASR for both benign scenarios is the same.
@@
(a) Scenario 1 (poor lighting)
(b) Scenario 2 (good lighting)
@@
TABLE VI: ASR in the real-world experiment.
@@ X Discussion
X-A Universal vs. Individual Attacks
@@
For each Emergency Landing frame we synthesize a scene-specific visual prompt and evaluate it on GPT-4o. These single-scene attacks raise the in-sample ASR to $84.35\%$ (vs. $68.10\%$ for the universal patch reported in Tab. IV), but they fail to generalize: applied to the held-out Transferability Images the single-scene ASR falls to $48.44\%$ (compared with $71\%$ for the universal attack reported in Tab. V). In short, view-specific (omniscient) attacks can substantially boost in-sample success, but their advantage collapses on novel scenes—highlighting why CHAI emphasizes robust, reusable prompts rather than per-view overfitting.
X-B Insights from Experiments
@@
LVLMs are susceptible to attacks in different languages: A drawback of English-only prompts is that English-speaking bystanders can easily read and flag the malicious sign. To gauge whether language switching thwarts human detection while still fooling the model, we repeat the experiment with Chinese and Spanish. This allows us to test cross-lingual robustness and explore attacks that are less conspicuous in an English-speaking environment.
In Fig. 12(a) and 12(b) we present an attack using Chinese and Spanish. For Spanish, we use the words Gire izquierda (Turn left) [^1] Gire izquierda is a short version of the grammatically correct string Gire a la izquierda. and an equivalent for Chinese. While humans who are not speakers of those languages may not understand, the LVLM can still understand the signs. In both cases, the LVLM commands the vehicle to turn left, explaining that turning left is appropriate to follow the traffic signal.
While those attacks may help to hide the adversarial actions, Chinese and Spanish speakers can still understand them. Fig. 12(b) presents an attack in Spanglish. That is, a combination of (Span)ish and Eng(lish). In particular, we found that the visual prompt Turn izquierda successfully induced the LVLM to output Turn left. Thus, we conclude that LVLMs are susceptible to visual prompt attacks using various languages, and an attacker can exploit that capability.
(a) Benign
(b) Prompt: Turn left
Figure 11: Example of an attack against DriveLM. We omit other DriveLM questions for simplicity.
@@
Closed-Loop Discussion: We implemented the emergency-landing task with the Microsoft AirSim UAV plugin. The virtual city includes dynamic lighting, textured façades, and camera shake, forcing the model to handle variations in viewpoint and illumination.
@@
We then attach a CHAI sign reading Safe to land to the hazardous roof and replay the identical flight trajectory. As illustrated in Fig. 14, the drone is diverted to the unsafe roof in $92\%$ of the runs, despite continuous changes in altitude, angle, and lighting, again showing the robustness of CHAI under realistic embodied AI scenarios.
(a) Benign
(b) Prompt: Safe to land
Figure 14: Example of runs in AirSim.
XI Conclusions and Future Work
This work introduces CHAI attacks, a new family of structured command-hijacking threats that embed natural-language prompts into visual scenes to mislead LVLMs controlling embodied AI systems. By coupling a dictionary-guided semantic search with cross-entropy optimization over perceptual features, we craft universal signs that flip high-level decisions in three representative agents: drone emergency landing, DriveLM driving, and CloudTrack tracking, with success rates up to $93\%$ while remaining practical to deploy. We also validate CHAI on a robotic vehicle and show that printed prompts can reliably bias LVLM decisions under real-world conditions. Experiments further demonstrate that CHAI generalizes across languages, weather, and viewpoints, exposing an attack surface unreachable by classical pixel-level adversarial patches. We also examined classical adversarial patches and white-box optimization: Appendix B details why patches are more challenging in our scenarios, and Appendix C outlines that white-box optimization yields no significant gains.
These findings highlight the need for defenses that reason over text and vision. Future directions include: (1) Filter-Based Defense: adding filters to the input image space or the output text space to recognize text inputs in the image or injected prompts in the output. A joint filter incorporating both the image and the text spaces would also be possible to defend against structured command hijacking. (2) Safety Alignment: fine-tuning and safeguarding LVLMs to defend directly against such visual prompts, such as discouraging the model from recognizing texts in the input space. Some prior works [43] have been done in the text-to-image model space, and the application to LVLMs is also largely unknown and would be an interesting direction to pursue. (3) Provable Defense: in the past, people have explored provable defenses [44] against patch attacks. The general idea would also be possible for visual prompts, because a patch may also occupy partially or entirely on the visual prompts, thus making a provable defense possible.
@@ Acknowledgments
We would like to thank Maciej Buszko for a proof of concept of these attacks for the advanced security class at UCSC in the winter quarter of 2025.
@@
Originality: We used an LLM (GPT-4o) to generate Fig. 1. Although we initially attempted to create this illustration manually, one of the authors experimented with a text prompt describing the intended design. The resulting image more effectively conveyed our idea than our manual sketches, so we adopted it for clarity of presentation.
@@
LVLMs selection: We selected GPT-4 to create the dictionary for our attack due to its superior capabilities. As target LVLMs, we utilize GPT and InterVL to demonstrate that our attack is applicable to both open-source and proprietary models.
@@
Computational resources: To run our experiments, we use queries to GPT through the API and a local computer to run InternVL. This computer features an Intel Core i9-13900K processor and an NVIDIA RTX A6000 graphics card.
Design decisions to decrease the use of LLMs: To ensure a fair use of the resources, we made several decisions that decrease the number of queries to VLMS. First, we constrained the number of iterations used to solve the optimization problem. Second, we reduced the number of decision variables, limiting it to color and the text of the visual prompt. Third, we limited the set of possible words to ten to decrease the search space. While we could increase the attack success by integrating more decision variables, a larger dictionary, or allowing the optimizer to run for longer, we would make the optimization problem more costly, implying the need for more queries.
Dataset creation: We collected data for DriveLM from Nuscenes [45], a public dataset. For the emergency landing and CloudTrack, we utilized a high-fidelity simulation, CARLA, which is a common tool to carry out in the field of autonomous vehicles and security. Finally, for the real-world experiment, we used our equipment, ensuring that no people were present during the experiment.

### v2

@@ VII Joint Semantic and Perceptual Optimization
The cross-entropy method defines a probability distribution in the search space $\Pi$, assigning a higher probability mass to regions that produce better objective values. Let $\Omega$ denote the (unknown) optimal distribution. Our goal is to approximate $\Omega$ by a parametric distribution $p_{\alpha}$, with parameters $\alpha\in P$, that minimizes the divergence $\mathcal{KL}(\Omega\,,\,p_{\alpha})$. In the ideal case, $\mathcal{KL}(\Omega,p_{\alpha})=0$.
@@
By using the first-order optimality conditions, we can get the solution to Equation (5) in a closed form depending on the form of $p_{\alpha}$ and $\Omega$. In particular, we consider a piecewise uniform distribution and divide the space $\Pi$ into $m$ disjoint subsets $C_{1},...,C_{m}\subset\Pi$. Then, the jthj-th element of $\alpha_{h}$, denoted as $\alpha_{h,j}$, represents the probability associated with the maximum being in $C_{j}$. We then update $\alpha$ as,
[…]
with $\gamma_{i}=\frac{\Omega(\pi_{i})}{p_{\alpha_{i}}(\pi_{i})}$. Refer to [37, 38] for the derivation.
@@ VIII-A Applications
We also implement a variation of DriveLM for experiments with real-world robotic vehicles, where we deploy the attack on a printed surface.
@@
Table III presents the optimization variables for each application. For most applications, we optimize the sign letter background colors in RGB space and the prompt.
Figure 6: Landing: Known (left) and Transferability (right).
Figure 7: DriveLM: Known (left) and Transferability (right).
GPT InternVL Application No Attack SceneTAP CHAI No Attack SceneTAP CHAI Landing $0.00$ p m 0.00 $6.25$ p m 3.19 72.75 7.15 $15.50$ p m 8.96 $26.5$ p m 12.99 66.43  9.89 CloudTrack $0.00$ p m 0.00 $77.00$ p m 8.01 92.00 4.10 $14.50$ p m 8.87 $42.50$ p m 14.25 92.50  9.10 DriveLM $6.22$ p m 2.93 $55.67$ p m 6.22 81.78 5.00 $29.41$ p m 7.09 $47.67$ p m 12.85 54.74 13.07
TABLE IV: Attack Success Rate (ASR) in the Known Images Datasets.
@@ VIII-B Evaluation Metrics
where $y_{i}^{t}\in\mathcal{Y}$ is the output that the attacker wants the LVLM to generate.
For this evaluation, we follow a similar approach to previous works [11], where we also evaluate the ASR in the scenario without an attack. This accounts for the stochasticity inherent to LVLMs; a query with the same image and prompt may create different outputs.
@@ VIII-C Datasets
For each application, we constructed two datasets: Known Images, used during attack optimization, and Transferability Images, held out to evaluate generalization (i.e., images that we do not show our optimizer). Figs. 6-7 illustrate this setup: each figure pairs an example from the Known Images set (left) with one from the Transferability Images set (right), highlighting the diverse conditions used to test CHAI on images unseen during optimization.
Details of the dataset construction, their diversity, and the labeling (including ASR errors in the benign case) can be found in Appendix A.
@@ VIII-D Implementation Details
LVLM: We test our attack against GPT-4o [40], a proprietary model, and InternVL2.5 8B [41], an open-weights model. We use commercial off-the-shelf models as several commercial controllers allow the deployment of those models to control the drone [42].
@@
Closed-loop implementation: We use AirSim [44], which is an open-source high-fidelity simulator for drones, to create a closed-loop implementation of the landing application.
@@ IX Results
In this section, we implement CHAI attacks for the applications described in Section VIII. We begin by evaluating CHAI on the Known Images and benchmarking against SceneTAP, then assessing transferability on the held-out Transferability Images. We then print physical signs of the optimized prompts and perform physical-world tests on a real robotic platform.
@@ IX-A Attack Success on Known Images
We implement CHAI and SceneTAP on Known Images. Table IV compares the ASR between both strategies. We draw the following conclusions.
@@
Application GPT InternVL No Attack CHAI No Attack CHAI Landing $0.00$ p m 0.00 71.38 6.34 $21.88$ p m 8.46 52.22 11.01 CloudTrack $0.00$ p m 0.00 91.00 7.18 $15.00$ p m 10.51 66.50  9.88 DriveLM $2.08$ p m 5.20 81.92 4.98 $33.08$ p m 15.90 51.25 15.12
TABLE V: ASR [%] in the Transferability Images Datasets.
@@ IX-B Attack Success on Transferability
We now study how our attack transfers to unknown scenarios. We apply the CHAI attacks that we obtain with the Known Images, apply them to the Transferability Images, and present the results in Table V. Notice that we cannot include SceneTAP in these results as SceneTAP is optimized only per image, and thus cannot generalize to previously unseen images.
@@
IX-C Real World Deployment
Attack setup: We position a second robotic vehicle as an obstacle directly ahead of the victim vehicle so that, under benign conditions, the LVLM issues a stop command. The attacker places a printed adversarial prompt on or near the obstacle and aims to cause the LVLM to output a proceed command instead; an attack succeeds if this induced command causes the victim to move forward and collide.
We consider two attack scenarios, which we illustrate in Figs. 8(b)-8(c):
@@ IX-C Real World Deployment
Table VI presents the ASR for both scenarios and LVLMs. When testing the case without an attack, we remove the objects where we place the CHAI signs. Consequently, the ASR for both benign scenarios is the same.
(a) Attack-vehicle (poor lighting)
(b) Off-vehicle (good lighting)
@@
TABLE VI: ASR [%] in the real-world experiment.
@@ X Discussion
X-A Ablation Study
We conducted an ablation study in Table VII considering two scenarios: No Dictionary: we select the word from SceneTAP when it does not generate the optimal visual prompt (landing and CloudTrack), or a random word from our dictionary. Then we optimize the perceptual features. No Perceptual: we perform the attack using each sentence in our dictionary with naive perceptual features (e.g., black and white colors). Since DriveLM has several possible outputs, we present results only when the attacker wants the car to proceed (excluding stop and turn commands). Consequently, we can understand the effects of each part of the attack.
Semantic Features Outweigh Perceptual Features: Semantic selection has a significant impact on attack success. Even after optimizing perceptual features, poor word selection results in a low ASR (CHAI-No Dictionary). Once the attacker finds a better word, the ASR increases, even when using non-optimized perceptual features (CHAI-No Perceptual).
Perceptual Features Further Increase the Attack Performance: Optimizing the perceptual features with the correct word further increases the ASR. In particular, for certain applications such as landing with GPT, the ASR increases from $25.50\%$ (without optimization) to $72.75\%$ (with optimization).
CHAI No Dictionary CHAI No Perceptual CHAI Full Landing GPT  6.30 25.50 72.75 InternVL 26.00 58.75 66.43 CloudTrack GPT  0.00 86.54 92.00 InternVL 40.52 84.85 92.50 DriveLM GPT 58.33 66.15 84.23 InternVL 43.13 50.62 55.21
TABLE VII: ASR [%] for the ablation study.
X-B Universal vs. Individual Attacks
@@ X-A Ablation Study
For each Emergency Landing frame, we synthesize a scene-specific visual prompt and evaluate it on GPT-4o. These single-scene attacks raise the in-sample ASR to $84.35\%$ (vs. $72.75\%$ for the universal patch reported in Table IV), but they fail to generalize: applied to the held-out Transferability Images, the single-scene ASR falls to $48.44\%$ (compared with $71\%$ for the universal attack reported in Table V). In short, view-specific (omniscient) attacks can substantially boost in-sample success, but their advantage collapses on novel scenes—highlighting why CHAI emphasizes robust, reusable prompts rather than per-view overfitting.
(a) Benign
(b) CHAI attack
Figure 10: Attack against DriveLM. We omit other DriveLM questions for simplicity.
X-C Insights from Experiments
@@
LVLMs are susceptible to attacks in different languages: A drawback of English-only prompts is that English-speaking bystanders can easily read and flag the malicious sign. To gage whether language switching thwarts human detection while still fooling the model, we repeated the experiment with Chinese and Spanish. This allows us to test cross-lingual robustness and explore attacks that are less conspicuous in an English-speaking environment.
In Fig. 11(a) and 11(b) we present an attack using Chinese and Spanish. For Spanish, we use the words and an equivalent for Chinese. While humans who are not speakers of those languages may not understand, the LVLM can still understand the signs. In both cases, the LVLM commands the vehicle to turn left, explaining that turning left is appropriate to follow the traffic signal.
While those attacks may help to hide the adversarial actions, Chinese and Spanish speakers can still understand them. Fig. 11(c) presents an attack in Spanglish. That is, a combination of (Span)ish and Eng(lish). In particular, we found that the visual prompt Turn izquierda successfully induced the LVLM to output Turn left. Thus, we conclude that LVLMs are susceptible to visual prompt attacks using various languages, and an attacker can exploit that capability.
@@
Closed-Loop Discussion: We implemented the emergency-landing task with the Microsoft AirSim UAV plugin [23]. The virtual city includes dynamic lighting, textured façades, and camera shake, forcing the model to handle variations in viewpoint and illumination.
@@
We then attach a CHAI sign reading Safe to land to the hazardous roof and replay the identical flight trajectory. As illustrated in Fig. 13(b), the drone is diverted to the unsafe roof in $92\%$ of the runs, despite continuous changes in altitude, angle, and lighting, again showing the robustness of CHAI under realistic embodied AI scenarios.
(a) Benign.
(b) CHAI attack.
Figure 13: Attacking a closed-loop landing scenario in AirSim.
X-D Future Work
Defenses: Our findings highlight the need for defenses that reason over text and vision. Future directions include: (1) Filter-Based Defense: adding filters to the input image space or the output text space to recognize text inputs in the image or injected prompts in the output. A joint filter incorporating both the image and the text spaces would also be possible to defend against structured command hijacking. (2) Safety Alignment: fine-tuning and safeguarding LVLMs to defend directly against such visual prompts, such as discouraging the model from recognizing texts in the input space. Some prior works [48] have been done in the text-to-image model space, and the application to LVLMs is also largely unknown and would be an interesting direction to pursue. (3) Provable Defense: in the past, people have explored provable defenses [49] against patch attacks. The general idea would also be possible for visual prompts, because a patch may also occupy partially or entirely on the visual prompts, thus making a provable defense possible.
@@ X-D Future Work
Domain-specific tuned LVLMs: In this paper, we studied attacks in Out of the Box models, as commercial autonomous vehicles allow their deployment [50]. While domain-specific training can improve LVLM performance on specific tasks, it does not inherently address adversarial vulnerabilities, typically requiring targeted adversarial training [51]. We leave the study of adversarial training and its effects on agent and attack performance for future work.
XI Conclusions
This work introduces CHAI attacks, a new family of structured command-hijacking threats that embed natural-language prompts into visual scenes to mislead LVLMs controlling embodied AI systems. By coupling a dictionary-guided semantic search with cross-entropy optimization over perceptual features, we craft universal signs that flip high-level decisions in three representative agents: drone emergency landing, DriveLM driving, and CloudTrack tracking, with success rates up to $93\%$ while remaining practical to deploy. We also validate CHAI on a robotic vehicle and show that printed prompts can reliably bias LVLM decisions under real-world conditions. Experiments further demonstrate that CHAI generalizes across languages, weather, and viewpoints, exposing an attack surface unreachable by classical pixel-level adversarial patches. We also examined classical adversarial patches and white-box optimization: Appendix C details why patches are more challenging in our scenarios, and Appendix D outlines that white-box optimization yields no significant gains.
Naïve defenses against CHAI—such as requiring authentication of external instructions or disabling multimodal perception altogether (e.g., via OCR or text-filtering)—are only effective in highly constrained deployment settings. Authentication presumes a closed and well-defined set of trusted entities (e.g., the members of a household interacting with a smart speaker, but not accepting the commands of anyone else), an assumption that breaks down in open environments such as public streets, or disaster zones.
Similarly, suppressing multimodal inputs undermines the very capabilities that motivate the use of LVLMs for embodied robotics: it would prevent legitimate, situationally necessary interactions with unanticipated humans and degrade task performance. For example, there are several videos online showing how autonomous vehicles struggle with flaggers attempting to communicate safety-critical information with signs [52].
In general, these examples illustrate a fundamental security gap posed by indirect prompt injection attacks in the physical environment. Security for embodied AI cannot be reduced to filtering prompts in current or future sensor modalities (e.g., LiDAR, radar, etc.). Instead, embodied agents must decide whether an inferred instruction is appropriate given their physical state, environment, and mission objectives.
@@ Acknowledgments
We thank Maciej Buszko for developing an early proof-of-concept of a CHAI attack as part of the Advanced Security course at UCSC early in 2025.
@@
Originality: We used an LLM (GPT-5.2) to generate Fig. 1. Although we initially attempted to create this illustration manually, one of the authors experimented with a text prompt describing the intended design. The resulting image more effectively conveyed our idea than our manual sketches, so we adopted it for clarity of presentation.
@@
LVLMs selection: We selected GPT-4 to create the dictionary for our attack due to its superior capabilities. As target LVLMs, we utilize GPT and InternVL to demonstrate that our attack is applicable to both open-source and proprietary models.
@@
Computational resources: To run our experiments, we use API queries to GPT and a local computer to run InternVL. This computer features an Intel Core i9-13900K processor and an NVIDIA RTX A6000 graphics card.
Design decisions to decrease the use of LLMs: To ensure fair use of resources, we made several decisions that reduced the number of queries to LVLMs. First, we constrained the number of iterations used to solve the optimization problem. Second, we reduced the number of decision variables, limiting it to color and the text of the visual prompt. Third, we limited the set of possible words to ten to decrease the search space. While we could increase attack success by integrating more decision variables, expanding the dictionary, or allowing the optimizer to run longer, we would make the optimization problem more costly, requiring more queries.
Dataset creation: We collected data for DriveLM from Nuscenes [53], a public dataset. For the emergency landing and CloudTrack, we used a high-fidelity simulation, CARLA, which is a common tool in the field of autonomous vehicles and security. For the emergency landing application, we used AirSim inside a city environment based on Unreal Engine 4. Finally, for the real-world experiment, we used our equipment, ensuring no people were present.